# Generic motives and motivic cohomology of fields

### v1

@@ 3.1. Generic motives and points for the homotopy $t$-structures
The family of exact functors:
$\operatorname{DM}(k)\rightarrow\mathrm{Mod}_{R},\mathbf{E}\mapsto\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM}(k)}(\operatorname{\hat{M}}(E)\{i\}[n],\operatorname{M}(X))$
indexed by a function field $E/k$ and a pair of integers $(n,i)\in\mathbb{Z}^{2}$, is conservative.
In other words, $\big{(}\operatorname{\hat{M}}(E)\{i\}\big{)}_{E/k,i\in\mathbb{Z}}$ is a conservative family of points for the homotopy $t$-structure, and this implies that the family of functors:
$\operatorname{HM}(k)\rightarrow\mathrm{Mod}_{R},F_{*}\mapsto\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{HM}(k)}\big{(}\operatorname{\underline{H}}_{0}(\operatorname{\hat{M}}(E))\{i\},F_{*}\big{)}=\hat{F}_{-i}(E)$
is conservative.
The corollary follows from the fact that the homotopy $t$-structure is non-degenerated and from the corresponding fact for homotopy sheaves: [Dég07, Prop. 4.7].
@@
Of course, one also obtains that $(\operatorname{\hat{M}}(E))_{E/k}$ is a conservative family of points for the homotopy $t$-structure on $\operatorname{DM^{eff}}(k)$ (resp. for homotopy sheaves).
The preceding proposition is the first step to compute morphisms in the homotopy category of generic motives. More precisely:
Lemma 3.1.11.
Let $E$, $F$ be function fields over $k$, and $n,m\in\mathbb{Z}$. Let $X$, $Y$ be respective smooth $k$-models of $E$ and $F$. Then there exists a canonical isomorphism, where $V$ and $U$ runs over non-empty open subsets:
$\pi_{0}\left(\underset{V\subset Y}{\operatorname{holim}}\ \underset{U\subset X}{\operatorname{hocolim}}\operatorname{Map}_{\operatorname{DM}(k)}(\operatorname{M}(U),\operatorname{M}(V))\right)$
$\qquad\simeq\lim_{V\subset Y}\underset{U\subset X}{\operatorname{colim}}\left(\pi_{0}\operatorname{Map}_{\operatorname{DM}(k)}(\operatorname{M}(U),\operatorname{M}(V))\right).$
Proof.
As filtered colimits are exact, one first deduces, for any $V\subset Y$ an isomorphism:
$\pi_{0}\left(\underset{U\subset X}{\operatorname{hocolim}}\operatorname{Map}_{\operatorname{DM}(k)}(\operatorname{M}(U),\operatorname{M}(V))\right)\simeq\underset{U\subset X}{\operatorname{colim}}\left(\pi_{0}\operatorname{Map}_{\operatorname{DM}(k)}(\operatorname{M}(U),\operatorname{M}(V))\right).$
According to the preceding proposition, one further gets:
$\underset{U\subset X}{\operatorname{colim}}\left(\pi_{0}\operatorname{Map}_{\operatorname{DM}(k)}(\operatorname{M}(U),\operatorname{M}(V))\right)\simeq\hat{\operatorname{\underline{H}}}_{0}(V)(E).$
But if $j:W\subset V$ is a non-empty open subset, it follows from [Dég07, Cor 3.22] that the morphism of homotopy invariant sheaves with transfers $\operatorname{\underline{H}}_{0}(W)\rightarrow\operatorname{\underline{H}}_{0}(V)$ is an epimorphism. This implies that the inductive sytem: $\left(\operatorname{\underline{H}}_{0}(V)(E)\right)_{V\subset Y}$ satisfies the Mittag-Leffler condition. In particular, one deduces (see e.g. [Goe96, Th. B]):
$\pi_{0}\left(\underset{V\subset Y}{\operatorname{holim}}\ \hat{\operatorname{\underline{H}}}_{0}(V)(E)\right)\simeq\lim_{V\subset Y}\hat{\operatorname{\underline{H}}}_{0}(V)(E)$
and this concludes. ∎
Remark 3.1.12.
Note in particular that one deduces a fully faithful functor:
$\operatorname{Ho}\operatorname{DM^{(0)}}(k)\rightarrow\operatorname{pro-\mskip-3.0mu}\operatorname{Ho}\operatorname{DM_{gm}}(k)$
whose image is the category of generic motives of [Dég08b, Def. 3.3.1].
Combining the above proposition and the preceding lemma, one deduces an embedding of the homotopy category of generic motives into an abelian category.
Theorem 3.1.13.
The functor
$\operatorname{Ho}\operatorname{DM^{(0)}}(k)\rightarrow\operatorname{pro-\mskip-3.0mu}\operatorname{HM}(k),\operatorname{\hat{M}}(E)\{n\}\mapsto\operatorname{\underline{H}}_{0,*+n}(\operatorname{\hat{M}}(E))=\operatorname{\underline{H}}_{0,*}(\operatorname{\hat{M}}(E))\{n\}$
is fully faithful.
3.2. Comparison with Beilinson’s approach
3.2.1.
Recall that the stable $\infty$-category $\operatorname{DM}(k)$ is generated by its compact objects, which are precisely the geometric motives (as recalled in 2.1.3). This readily implies an interpretation of the “big” category of motives $\operatorname{DM}(k)$ as ind-geometric objects:
Thus using that, after inverting the characteristic exponent $e$ of $k$, $\operatorname{DM_{gm}}(k)_{e}$ is made of rigid objects (see loc. cit.), one further deduces the following interpretation of pro-geometric motives.
Proposition 3.2.2.
The canonical functor:
$\big{(}\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k)_{e}\big{)}^{op}\rightarrow\operatorname{DM}(k)_{e},(\mathbf{E}\simeq\underset{i\in I}{\text{\rm``lim''}}\,\mathbf{E}_{i})\mapsto\operatorname{hocolim}_{i\in I}\mathbf{E}_{i}^{\vee}$
is an equivalence of categories.
3.2.3.
The preceding functor maps the motive $\operatorname{M}(X)$ of a $k$-scheme of finite type $X$ to its dual $\operatorname{h}(X):=\operatorname{\underline{Hom}}(\operatorname{M}(X),\mathbbm{1})=f_{*}(\mathbbm{1}_{X})$, the cohomological motive of $X$ (see 2.1.3).
The preceding proposition allows to translate the results of 3.1.3 to cohomological motives (without necessarily inverting $e$). There exists a right Kan extension $\operatorname{\hat{h}}$ of the functor $\operatorname{h}:\mathrm{Sm}_{k}\rightarrow\operatorname{DM}(k)$ along the inclusion $\nu:\mathrm{Sm}_{k}\rightarrow\overline{\mathrm{Sm}}_{k}$:
${\mathrm{Sm}_{k}^{op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu}$$\scriptstyle{\operatorname{h}}$${\operatorname{DM}(k)}$${\overline{\mathrm{Sm}}_{k}^{op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{\hat{h}}}$
Moreover, this extension can be computed as
$\operatorname{\hat{h}}(\mathcal{X})=\operatorname{hocolim}_{i\in I}\operatorname{h}(X_{i})$
for an essentially smooth scheme $\mathcal{X}$ with a pro-presentation $\mathcal{X}=\lim_{i\in I}X_{i}$.
In particular, one can defined the cohomological motive attached to function fields $E/k$:
$\operatorname{\hat{h}}(E)=\underset{U\subset X}{\operatorname{hocolim}}\operatorname{h}(U)$
for any smooth model $X$ of $E/k$. This is exactly Beilinson’s “motive of the generic point” of $X$ from [Bei02, 3.1] — after identifying Voevodsky’s motivic spectra $\operatorname{DM}(k)$ with ind-geometric motives according to (3.2.1.a). The preceding proposition readily gives the following comparison with the preceding notion of generic motives.
Proposition 3.2.4.
The canonical functor
$\operatorname{DM^{(0)}}(k)_{e}^{op}\rightarrow\operatorname{DM}(k)_{e},\operatorname{\hat{M}}(E)\{i\}\mapsto\operatorname{\hat{h}}(E)\{-i\}$
is fully faithful and induces an anti-equivalence of categories between the category of generic motives and the sub-category of motivic spectra made by $\mathbb{G}_{m}$-twists of Beilinson’s cohomological motives of generic points.
Beilinson has already explained the conjectural injective nature of the ind-motives $\operatorname{\hat{h}}(E)$ with respect to the motivic $t$-structure (see [Bei02, Cor. 3.3]). In fact, if one uses the homotopy $t$-structure, one can establish analogous but non-conjectural facts for generic motives. Let us first recall some definitions (see [Bei02, before Cor. 3.3], [GLVdB21, §6.1]).
Definition 3.2.5.
Let $\mathscr{T}$ be a $t$-category. [^20] A triangulated category equipped with a $t$-structure. An object $M$ of the heart $\mathscr{T}^{\heartsuit}$ is called $t$-projective if $\operatorname{Hom}_{\mathscr{T}}(M,\mathscr{T}_{>0})=0$.
An object $K$ of $\mathscr{T}$ is called derived $t$-projective if the following two conditions hold:
@@ 3.2. Comparison with Beilinson’s approach
One easily derives that $t$-projective for an object of the heart of $\mathscr{T}$ implies derived $t$-projective. Also, a derived $t$-projective object is necessarily (homologically) non-negative.

### v3

@@ 3.1. Generic motives and points for the homotopy $t$-structures
For any function field $E/k$ and integer $i\in\mathbb{Z}$, the $\infty$-functor
$\operatorname{\mathscr{D}M}(k)\rightarrow\mathscr{S}p,\mathbf{E}\mapsto\operatorname{Map}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M}(k)}(\operatorname{\hat{M}}(E)\{i\},\mathbf{E})$
is $t$-exact, with respect to the homotopy $t$-structure on the left and the canonical $t$-structure on the right. Moreover, the family of these $t$-exact $\infty$-functors indexed by a function field $E/k$ and an integer $i\in\mathbb{Z}$, is conservative.242424i.e. a morphism in $\operatorname{\mathscr{D}M}(k)$ is an isomorphism if and only if its image by all the functors of the family considered is an isomorphism. See also 6.2.1.
This corollary is a compact form of the following two assertions:
  • A motivic spectrum $\mathbf{E}$ is non-negative (resp. negative) for the homotopy $t$-structure if and only if for any $(E,i)$ as above, the spectrum $\operatorname{Map}(\operatorname{\hat{M}}(E)\{i\},\mathbf{E})$ is non-negative (resp. negative).
  • A morphism of motivic spectra $f:\mathbf{E}\rightarrow\mathbf{F}$ is an isomorphism if and only if for any $(E,i)$ as above, the map $\operatorname{Map}(\operatorname{\hat{M}}(E)\{i\},f)$ is an isomorphism in the $\infty$-category of spectra.
Proof.
Indeed, 3.1.7 applied to the case $\mathcal{O}=E$ gives the formula:
$\pi_{n}\operatorname{Map}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M}(k)}(\operatorname{\hat{M}}(E)\{i\},\mathbf{E})=\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M}(k)}(\operatorname{\hat{M}}(\mathcal{O})\{i\}[n],\mathbf{E})=\hat{\operatorname{\underline{H}}}_{n,-i}(\mathbf{E})(\mathcal{O}).$
Then the corollary follows from the definition of the homotopy $t$-structure in 2.1.4, and one of the main results of Voevodsky’s theory of motivic complexes: the fact that function fields form a conservative family for homotopy sheaves (see [Dég07, Prop. 4.7]). ∎
@@
In more classical terms, one gets that the family of exact functors with values in the category of abelian groups:
(3.1.11.a) $\operatorname{HM}(k)\rightarrow\mathrm{Ab},F_{*}\mapsto\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{HM}(k)}\big{(}\operatorname{\underline{H}}_{0}(\operatorname{\hat{M}}(E))\{i\},F_{*}\big{)}=\hat{F}_{-i}(E)$
is conservative. In fact, this assertion can be deduced directly from Voevodsky’s result [Dég07, Prop. 4.7].
3.1.12.
Beilinson has already discussed the conjectural projective nature of generic motives with respect to the conjectural motivic $t$-structure (see [Bei02, Cor. 3.3]).
We will now formulate a similar, non-conjectural, property of generic motives with respect to the homotopy $t$-structure. Let us first recall from [GLVdB21, §6.1] that, given an arbitrary stable $\infty$-category $\mathscr{T}$ equipped with a $t$-structure, an object $K$ of $\mathscr{T}$ is called derived $t$-projective if the following two conditions hold:
@@
To address the case of generic motives, one needs an appropriate reformulation.
Definition 3.1.13.
Let $\mathscr{T}$ be an $\infty$-category equipped with a $t$-structure (see 6.4.10), and let $\mathcal{K}$ be a pro-object of $\mathscr{T}$. The canonical homological $\infty$-functor $H_{0}:\mathscr{T}\rightarrow\mathscr{T}^{\heartsuit}$, [^25] According to our usual abuse of notation, the target is really the nerve of the category $\mathscr{T}^{\heartsuit}$. Of course, this $\infty$-functor is equivalent to the more classical one: $H_{0}:\operatorname{Ho}(\mathscr{T})\rightarrow\operatorname{Ho}(\mathscr{T})^{\heartsuit}$, using 6.1.7(2) and the identification $\mathscr{T}^{\heartsuit}=\operatorname{Ho}(\mathscr{T})^{\heartsuit}$. induces a functor on pro-categories
$\operatorname{pro-\mskip-3.0mu}\mathscr{T}\rightarrow\operatorname{pro-\mskip-3.0mu}{\mathscr{T}^{\heartsuit}}$
that we will abusively denote by $H_{0}$. [^26] Note that according to Section 6.6, the right-hand side is the ordinary category of pro-objects of $\mathscr{T}^{\heartsuit}$, which is therefore abelian. However, we will not use this fact in the sequel.
One says that $\mathcal{K}$ is derived pro-$t$-projective if the following conditions hold:
  1. (i)
    For all $\mathbf{E}$ in $\mathscr{T}$, $\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\mathscr{T}}(\mathcal{K},\mathbf{E})\simeq\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\mathscr{T}^{\heartsuit}}(H_{0}(\mathcal{K}),H_{0}(\mathbf{E}))$.
  2. (ii)
    The functor $\mathscr{T}^{\heartsuit}\rightarrow\mathrm{A}b,A\mapsto\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\mathscr{T}^{\heartsuit}}(H_{0}(\mathcal{K}),A)$ is exact.
The second condition sometimes appears in the literature under the name “pro-projective”. The next proposition gives us the desired property of generic motives.
Proposition 3.1.14.
Let $E/k$ be a function field and $i\in\mathbb{Z}$ an integer. Then the generic motive $\operatorname{\hat{M}}(E)\{i\}$ is derived pro-$t$-projective with respect to the homotopy $t$-structure on $\operatorname{\mathscr{D}M}(k)$.
Proof.
We prove that the pro-motive $\operatorname{\hat{M}}(E)\{i\}$ satisfies the conditions of the preceding definition. In fact, point (i) follows from 3.1.7, while point (ii) follows from 3.1.11. ∎
3.2. Categorical variants and presentations of generic motives
In this section, we will show how one can compare various constructions of generic motives. We start with a key lemma.
Lemma 3.2.1.
Let $E$, $F$ be function fields over $k$, and $n\in\mathbb{Z}$. Let $X$, $Y$ be smooth $k$-models of $E$ and $F$, respectively. Then there exists a canonical isomorphism, where $V$ and $U$ range over non-empty open subsets:
$\pi_{0}\left(\underset{V\subset Y}{\operatorname{lim}}\ \underset{U\subset X}{\operatorname{colim}}\operatorname{Map}_{\operatorname{\mathscr{D}M}(k)}(\operatorname{M}(U),\operatorname{M}(V)\{n\})\right)$
$\qquad\simeq\underset{V\subset Y}{\operatorname{lim}}\underset{U\subset X}{\operatorname{colim}}\left(\pi_{0}\operatorname{Map}_{\operatorname{\mathscr{D}M}(k)}(\operatorname{M}(U),\operatorname{M}(V)\{n\})\right)$
where we emphasize that on the left-hand side the colimit and limit are taken within the $\infty$-category of spaces, while on the right-hand side, they are taken in ordinary categorical sense, in the category of abelian groups.
Proof.
Since filtered colimits are exact, one first deduces the following isomorphism for any $V\subset Y$:
$\pi_{0}\left(\underset{U\subset X}{\operatorname{colim}}\operatorname{Map}_{\operatorname{\mathscr{D}M}(k)}(\operatorname{M}(U),\operatorname{M}(V)\{n\})\right)\simeq\underset{U\subset X}{\operatorname{colim}}\left(\pi_{0}\operatorname{Map}_{\operatorname{\mathscr{D}M}(k)}(\operatorname{M}(U),\operatorname{M}(V)\{n\})\right).$
Applying 3.1.7, one further gets:
$\underset{U\subset X}{\operatorname{colim}}\left(\pi_{0}\operatorname{Map}_{\operatorname{\mathscr{D}M}(k)}(\operatorname{M}(U),\operatorname{M}(V)\{n\})\right)\simeq\hat{\operatorname{\underline{H}}}_{0,n}(V)(E).$
But if $j:W\subset V$ is a non-empty open subset, it follows from [Dég07, Cor 3.22] that the morphism of homotopy invariant sheaves with transfers $\operatorname{\underline{H}}_{0}(W)\rightarrow\operatorname{\underline{H}}_{0}(V)$ is an epimorphism. Voevodsky’s cancellation theorem implies that the same is true for $\operatorname{\underline{H}}_{0,n}(W)(E)\rightarrow\operatorname{\underline{H}}_{0,n}(V)(E)$. [^27] We need the cancellation theorem only in the case $n>0$ as Voevodsky’s $(-1)$-functor — see (2.1.4.a) — is exact. This implies that the inductive system: $\left(\operatorname{\underline{H}}_{0,n}(V)(E)\right)_{V\subset Y}$ satisfies the Mittag-Leffler condition. In particular, one deduces (see e.g. [Goe96, Th. B]):
$\pi_{0}\left(\underset{V\subset Y}{\operatorname{lim}}\ \hat{\operatorname{\underline{H}}}_{0,n}(V)(E)\right)\simeq\underset{V\subset Y}{\operatorname{lim}}\left(\hat{\operatorname{\underline{H}}}_{0,n}(V)(E)\right)$
and this concludes the proof. ∎
3.2.2.
For the next statement, we consider the following composition of $\infty$-functors:
$\operatorname{\mathscr{D}M^{(0)}}(k)\hookrightarrow\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k)\rightarrow\operatorname{pro-\mskip-3.0mu}{\left(\operatorname{Ho}\operatorname{\mathscr{D}M_{gm}}(k)\right)}=\operatorname{pro-\mskip-3.0mu}{\operatorname{DM_{gm}}(k)}$
where the last map is obtained by applying the pro-object construction to the natural map $\operatorname{\mathscr{D}M_{gm}}(k)\rightarrow\operatorname{Ho}\operatorname{\mathscr{D}M_{gm}}(k)$. It follows from abstract non-sense (6.1.7(2)) that this defines an (ordinary) exact functor:
(3.2.2.a) $\operatorname{Ho}\operatorname{\mathscr{D}M^{(0)}}(k)\rightarrow\operatorname{pro-\mskip-3.0mu}{\operatorname{DM_{gm}}(k)}.$
We now deduce the following result, which allows us to compare 3.1.4 with the older definition of generic motives used in [Dég08b].
Proposition 3.2.3.
The functor (3.2.2.a) is fully faithful, with essential image given by the ordinary category of generic motives defined in [Dég08b, Def. 3.3.1].
In other words, the ordinary category of generic motives defined in loc. cit. is equivalent to the homotopy category of the $\infty$-category of generic motives $\operatorname{\mathscr{D}M^{(0)}}(k)$ defined here.
Proof.
One should be careful that this is not a tautology. Indeed, upon unwinding definitions and using Formula (6.6.4.a), we find that this is precisely the content of the preceding lemma. ∎
We also remark that, by combining the preceding result with 3.1.7, one can compute homotopy classes of morphisms of generic motives inside an abelian category, as expressed in the following proposition. [^28] For ordinary generic motives, this was already remarked in [Dég08b, Cor. 3.4.7].
Proposition 3.2.4.
The functor
$\operatorname{Ho}\operatorname{\mathscr{D}M^{(0)}}(k)\rightarrow\operatorname{pro-\mskip-3.0mu}\operatorname{HM}(k),\operatorname{\hat{M}}(E)\{n\}\mapsto\operatorname{\underline{H}}_{0,*+n}(\operatorname{\hat{M}}(E))=:\operatorname{\underline{H}}_{0,*}(\operatorname{\hat{M}}(E))\{n\}$
is fully faithful.
3.2.5.
We conclude this subsection by showing that, at the cost of inverting the characteristic exponent $e$ of the base field $k$, it is possible to consider ind-(cohomological motives) for modeling generic motives.
Recall that the presentable stable $\infty$-category $\operatorname{DM}(k)$ is generated by its compact objects, which are precisely the geometric motives (as recalled in 2.1.5). This observation leads to an interpretation of the “big” category of motives $\operatorname{DM}(k)$ as ind-geometric objects: [^29] This is well-known, but see [Lur09, Proposition 5.3.5.11] for a very general proof.
Working within a sufficiently large universe, we formally deduce the following equivalence of $\infty$-categories (see 6.6.2):
$(\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k))^{\mathrm{op}}=\operatorname{ind-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k)=\operatorname{\mathscr{D}M}(k).$
Thus, in principle, one can view generic motives as certain motivic spectra, provided one works in the opposite category. We can make this perspective more concrete by inverting the characteristic exponent $e$ of $k$. Indeed, as recalled in 2.1.5, every object of $\operatorname{DM_{gm}}(k)[1/e]$ is rigid, and for any smooth $k$-scheme $X$, the dual of $M(X)$ can be computed as the cohomological motive $h(X)$. Based on these observations, we obtain the following proposition:
Proposition 3.2.6.
The canonical (contravariant) $\infty$-functor:
$\big{(}\operatorname{\mathscr{D}M^{eff}_{gm}}(k)[1/e]\big{)}^{op}\rightarrow\operatorname{\mathscr{D}M}(k)[1/e],\big{(}\operatorname{\hat{M}}(E)\{i\}\big{)}\mapsto\underset{X\in\mathcal{M}(E)^{op}}{\operatorname{colim}}\big{(}\operatorname{h}(X)\{-i\}\big{)}$
where $X=\operatorname{Spec}(A)$ ranges over the smooth affine models of $E$ (3.1.2(2)) and the colimit is taken in the $\infty$-category $\operatorname{\mathscr{D}M}(k)$, is fully faithful.
Proof.
The proof reduces to the following computation of mapping spaces:
$\operatorname{Map}\big{(}\operatorname{colim}_{Y}\operatorname{h}(Y)\{-i\},\operatorname{colim}_{X}\operatorname{h}(X)\big{)}$ $\simeq\operatorname{lim}_{Y}\operatorname{Map}\big{(}\operatorname{h}(Y)\{-i\},\operatorname{colim}_{X}\operatorname{h}(X)\big{)}$
$\simeq\operatorname{lim}_{Y}\operatorname{colim}_{X}\operatorname{Map}\big{(}\operatorname{h}(Y)\{-i\},\operatorname{h}(X)\big{)}$
$\simeq\operatorname{lim}_{Y}\operatorname{colim}_{X}\operatorname{Map}\big{(}\operatorname{M}(X),\operatorname{M}(Y)\{i\}\big{)}$
$\simeq\operatorname{Map}\big{(}\operatorname{\hat{M}}(E),\operatorname{\hat{M}}(F)\{i\}\big{)}$
where $X$ and $Y$ range over the smooth affine models of some function fields $E$ and $F$, respectively. The first isomorphism is formal, the second one follows as $\operatorname{h}(Y)\{-i\}$ is compact, the third one follows by duality, and the last one follows from (6.6.4.a). ∎
3.2.7.
Ind-Chow motives as (co)models for generic motives. The preceding proposition shows that, by working in the opposite category and inverting the characteristic exponent $e$, one can replace generic motives by the following motivic spectrum: [^30] In fact, one can reproduce, in a dual fashion, the definitions of the preceding section. In particular, there exists a right Kan extension $\operatorname{\hat{h}}$ of the functor $\operatorname{h}:\mathrm{Sm}_{k}\rightarrow\operatorname{DM}(k)$ along the inclusion $\nu:\mathrm{Sm}_{k}\rightarrow\overline{\mathrm{Sm}}_{k}$: ${\mathrm{Sm}_{k}^{op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu}$$\scriptstyle{\operatorname{h}}$${\operatorname{DM}(k)}$${\overline{\mathrm{Sm}}_{k}^{op}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{\hat{h}}}$ such that $\operatorname{\hat{h}}(\mathcal{X})=\operatorname{hocolim}_{i\in I}\operatorname{h}(X_{i})$.
$\operatorname{\hat{h}}(E)\{i\}=\operatorname{colim}_{X\in\mathcal{M}(E)^{op}}\operatorname{h}(X)\{i\}$
where the colimit is computed in the $\infty$-category $\operatorname{\mathscr{D}M}(k)[1/e]$.
The sub-$\infty$-category of $\operatorname{\mathscr{D}M}(k)[1/e]$ is then equivalent, according to the preceding proposition, to the opposite of the $\infty$-category of generic motives after inverting $e$.